Find two positive integers whose product is 100 and whose sum is minimum


1
Expert's answer
2021-05-17T18:50:03-0400

Let the two positive numbers be x and y.


So, xy = 100 ...................Equation(1)


and x + y = S (To be minimized)


From Equation (1) ,


y = 100/x ..................Equation(2)

x + 100/x = S

(x2 + 100) / x = S

For critical points we let dS/dx = 0

dS/dx = [ 2x2 - 1( x2 + 100) ] / x2

dS/dx = ( x2 - 100 ) / x2 .....................Equation(3)


For critical points,

 ( x2 - 100 ) / x2 = 0

x = 10, -10


Since only positive values of x and y are needed, so we discard the negative value


Hence, x = 10


For the minimum value of sum d2S / dx2 > 0 at x = 10.

So differentiating Equation (3) with respect to x, we have


d2S / dx2 = [ 2x3 - 2x( x2 - 100 ) ] / x4


d2S / dx2 = 100 / x4


At x = 10, d2S / dx2 = 0.01 > 0


x= 10

y = 100/10 = 10


Hence, for the minimum value of sum the values of x and y are equal to 10 and 10 respectively.